Mr mobile legends want to take a table for his friends which has an area of 10 square meters. the length of the table has to 3 meters longer than the width. find the width and length of the table

Found 3 solutions by dkppathak, Theo, MathTherapy:
Answer by dkppathak(439)

Mr mobile legends want to take a table for his friends which has an area of 10 square meters. the length of the table has to 3 meters longer than the width. find the width and length of the table
solution
Let width is x mtrs
length will be x+3
area=lenghtxwidth
x(x+3)=10
x^2+3x-10=0
(x+5)(x-2)=0
x=-5 or x=2 width can not be negative
so x=2
length will be 5 mtrs

Answer by Theo(13342)

the area of the table top is 10 square meters.
the length of the table is 3 meters longer than the width.

let x = the length
let y = the width.
the area of the table top is equal to x * y

since the length of the table top is 3 meters longer than the width, you get:
x = 3y

area = x * y becomes area = 3y * y, after you replace x with 3y.

simplify to get:
area = 3y^2

since the area is 10 square meters, you get:
10 = 3y^2

divide both sides of this equation by 3 to get:
10/3 = y^2

solve for y to get:
y = plus or minus sqrt(10/3)

since y can't be negative, y must be equal to sqrt(10/3)
since x = 3y, then x must be equal to 3 * sqrt(10/3)

the area of the table is equal to x * y.
x * y = 3 * sqrt(10/3) * sqrt(10/3)
since sqrt(10/3) * sqrt(10/3) = 10/3, the area of the table becomes:
area = 3 * 10/3
simplify to get:
area = 10

this confirms the solution is correct.
the solution is:

length = 3 * sqrt(10/3)
width = sqrt(10/3)


in decimal format, .....

length = 5.477225575 meters.
width = 1.825741858 meters.




Answer by MathTherapy(10552)


Mr mobile legends want to take a table for his friends which has an area of 10 square meters. the length of the table has to 3 meters longer than the width. find the width and length of the table
The width and length are simply: 2 m, and 5 m, respectively.
That's ALL!! NOTHING more to this!